OpenStudy (anonymous):

evaluate log base 6 of the fourth root of 1/36

OpenStudy (wolf1728):

1/36 = (0.0277777778) (0.0277777778)^.25 0.4082482905 log base 6 (0.4082482905) = -.5 logarithm calculator located here: http://www.1728.org/logrithm.htm

OpenStudy (ranga):

1/36 = 1/6^2 = 6^(-2) fourth root of 1/36 = (1/36)^(1/4) = ( 6^-2 ) (1/4) = 6^(-2/4) = 6^(-1/2) log base 6 of 6^(-1/2) = -1/2